# ChunkRAG: Novel LLM-Chunk Filtering Method for RAG Systems

### v1

@@ Abstract
Ritvik Aggarwal           Ishneet Sukhvinder Singh           Ibrahim Allahverdiyev           Muhammad Taha Aslihan Akalin           Kevin Zhu Algoverse AI Research asli@algoverse.us, kevin@algoverse.us
@@ 3 Methodology
The primary aim of this work is to reduce the irrelevance and hallucinations in the responses generated by RAG systems by implementing a novel, fine-grained filtering process that evaluates the relevance of each chunk of retrieved information before it is used in the response generation phase.
@@
Semantic Chunking
The initial step in the methodology involves breaking down the document into semantically coherent chunks, referred to as semantic chunking. This approach ensures that each chunk contains closely related information, improving the relevance of retrieval operations.
Sentence Tokenization:
Each document is split into sentences using NLTK’s sent_tokenize function. This tokenization step enables the text to be processed at a more granular level.
Sentence Embeddings Generation:
Sentence embeddings are generated using OpenAI’s ‘text-embedding-3-small‘ model. Each sentence is transformed into a numerical vector representation, capturing its semantic meaning.
Similarity Calculation:
Cosine similarity is calculated between the embeddings of consecutive sentences. The similarity scores help determine how semantically related consecutive sentences are. If the similarity score drops below a predefined threshold (e.g., 0.7), it indicates a shift in topic.
Chunk Boundary Identification:
Based on similarity scores, chunk boundaries are identified at points where the similarity drops below the threshold, suggesting a change in topic.
Chunk Creation:
Sentences are grouped into chunks based on the identified boundaries. Each chunk is kept within a maximum length (e.g., 500 characters) to ensure semantic coherence and manageability.
Vector Store Creation
To facilitate efficient similarity-based retrieval, chunk embeddings are stored in a vector store:
Embedding Chunks:
@@
Each chunk is embedded using the same embedding model (‘text-embedding-3-small‘), resulting in numerical vector representations of the chunks.
Storing in Vector Store:
The embeddings are stored in a Chroma vector store, which allows for efficient similarity search operations during retrieval.
Retriever Initialization:
@@
A retriever is initialized to perform similarity-based searches. This retriever compares the embeddings of a query with those of the stored chunks to find the most relevant chunks.
Query Rewriting
To enhance retrieval performance, the user’s query is rewritten without changing its original intent:
LLM-Based Rewriting:
A Language Model (LLM), such as GPT-4o-mini, is used to rewrite the query. A prompt guides the LLM to produce an improved version of the query that enhances retrieval effectiveness.
Initial Filtering Method
To improve efficiency and ensure relevance, initial filtering methods are applied before computationally expensive scoring operations:
Remove Redundant Chunks:
Duplicate or highly similar chunks are eliminated using TF-IDF vectorization and cosine similarity measures to reduce redundancy.
Semantic Similarity Sorting:
The remaining chunks are sorted based on their similarity to the query, prioritizing more relevant chunks for subsequent processing.
Advanced LLM-Based Relevance Scoring
To assess the relevance of each chunk more accurately:
Initial LLM Scoring:
@@
The primary LLM assigns an initial relevance score to each chunk using a predefined prompt.
Self-Reflection:
The LLM reflects on its initial score and adjusts it if necessary to catch potential misjudgments.
Critic LLM Scoring:
A second LLM, termed as a critic, independently evaluates each chunk and assigns its own score to validate the initial assessment.
Final Score Calculation:
The final relevance score is obtained by averaging the initial, reflected, and critic scores to enhance accuracy and reliability.
Dynamic Threshold Determination
Instead of using a fixed threshold, a dynamic threshold is determined:
LLM-Based Threshold Determination:
@@
The LLM analyzes the distribution of final scores and suggests an optimal threshold for chunk relevance.
Threshold Application:
@@
Chunks with scores above the threshold are retained for answer generation.
Hybrid Retrieval and Re-Ranking
@@
BM25 Retriever: In parallel, a BM25 retriever is implemented to capture keyword-based retrieval. BM25 retriever is combined using an ensemble approach with equal weights (0.5 each).
Re-ranking Chunks:
@@
Cohere’s reranking model (‘rerank-english-v3.0‘) is applied to rank the chunks with relevance to deal with Lost in the middle problem.
Answer Generation
To generate the final answer for the user’s query:
Context Compilation:
The filtered chunks are compiled to form the context used by the LLM.
Answer Generation with LLM:
The LLM generates a concise answer based on the compiled context and the user’s original question using a prompt that ensures the answer is based solely on the provided context.
Evaluation
The system’s performance is evaluated using the accuracy metrics. The response generated by the LLM using the retriever is compared with the ground-truth (predetermined correct answer).
The comparison includes both exact and semantically correct answers. An evaluation LLM compares ground-truth and the generated answer and flags it either correct or incorrect.
The accuracy is calculated using:
@@ 4.2.2 Baselines With Retrieval
Advanced Retrieval-Augmented Generation: To benchmark ChunkRAG against more sophisticated RAG-based methods, we included advanced models that incorporate additional strategies to enhance performance. SAIL Luo et al. (2023) enhances standard RAG by instruction-tuning the language model on Alpaca instruction-tuning data, inserting top retrieved documents before the instructions to provide contextual information. Self-RAG Asai et al. (2024) further refines RAG by incorporating reflection tokens labeled by GPT-4 within the instruction-tuning data, enabling the model to better utilize retrieved information. Additionally, we considered CRAG Your et al. (2024), a recent approach that augments standard RAG with corrective strategies to improve retrieval quality by addressing low-quality retrieval results. CRAG serves as a direct comparison to our proposed ChunkRAG, highlighting the effectiveness of our chunk filtering mechanism in enhancing retrieval-augmented generation. Furthermore, we evaluated retrieval-augmented baselines with private data, including Ret-ChatGPT and RetLLaMA-chat, which integrate retrieval mechanisms with ChatGPT and the conversational variant of LLaMA2, respectively. We also assessed Perplexity.ai, an InstructGPT-based production search system that leverages retrieval-augmented techniques for generating accurate and relevant responses.

### v2

@@ Abstract
Ishneet Sukhvinder Singh           Ritvik Aggarwal           Ibrahim Allahverdiyev           Muhammad Taha Aslihan Akalin           Kevin Zhu          Sean O’Brien Algoverse AI Research asli@algoverse.us, kevin@algoverse.us, seobrien@ucsd.edu
@@ 3 Methodology
The primary aim of this work is to reduce irrelevance and hallucinations in the responses generated by Retrieval-Augmented Generation (RAG) systems through a novel, fine-grained filtering process that evaluates the relevance of each chunk of retrieved information before it is used in the response generation phase. Below, we describe the steps involved in a detailed, precise, and reproducible manner.
@@
Step 1: Semantic Chunking
  1. 1.
    Input Preparation: Start with a document $D$ and break it down into sentences using a sentence tokenizer (NLTK’s $sent\_tokenize$ function).
  2. 2.
    Embedding Generation: Use a pre-trained embedding model (e.g., text-embedding-3-small) to create a vector representation for each sentence.
  3. 3.
    Chunk Creation: Calculate the similarity between consecutive sentences using cosine similarity. If similarity drops below a threshold ($\theta=0.7$), mark a new chunk boundary. Group sentences accordingly, ensuring each chunk is under 500 characters.
Step 2: Vector Store Creation
  1. 1.
@@
    Embedding the Chunks: Convert each chunk into an embedding using the same model as above.
  2. 2.
    Storing Embeddings: Store all chunk embeddings in a vector database for easy retrieval during query matching.
Step 3: Retriever Initialization
  1. 1.
@@
    Setup Retriever: Initialize the retriever to compare incoming query embeddings with stored chunk embeddings to find the most relevant chunks.
Step 4: Query Rewriting
  1. 1.
    Query Enhancement: Rewrite the user’s original query using a language model (e.g., GPT-4omini) to make it more suitable for retrieval.
Step 5: Initial Filtering
  1. 1.
    Duplicate Removal: Use TF-IDF and cosine similarity to eliminate duplicate or overly similar chunks (similarity $>0.9$).
  2. 2.
    Sort by Relevance: Sort the remaining chunks based on similarity to the rewritten query.
Step 6: Advanced Relevance Scoring
  1. 1.
@@
    Initial Scoring: Assign a relevance score to each chunk using an LLM.
  2. 2.
    Score Refinement: Use self-reflection and a critic model to refine the initial scores, then calculate an average to obtain the final score.
Step 7: Thresholding for Relevance
  1. 1.
@@
    Set Dynamic Threshold: LLM analyzes the distribution of final scores and suggests an optimal threshold for chunk relevance.
  2. 2.
@@
    Filter Chunks: Retain only chunks with scores above the threshold.
Step 8: Hybrid Retrieval and Re-Ranking
  1. 1.
@@
    Dual Retrieval Methods: Combine BM25 and LLM-based retrieval methods to ensure a balanced retrieval strategy. BM25 retriever is implemented to capture keyword-based retrieval. BM25 retriever is combined using an ensemble approach with equal weights (0.5 each).
  2. 2.
@@
    Re-Rank Chunks: Cohere’s reranking model (‘rerank-englishv3.0‘) is applied to rank the chunks with relevance to deal with Lost in the middle problem
Step 9: Answer Generation
  1. 1.
    Compile Context: Collect the most relevant chunks to use as context.
  2. 2.
    Generate Response: Use an LLM to generate an answer based on the context, ensuring that only the information from retrieved chunks is used.
Step 10: Evaluation
  1. 1.
    Accuracy Calculation: Compare the generated answers against a set of correct answers to evaluate performance:
    Accuracy=Number of Correct AnswersTotal Number of $\text{Accuracy}=\frac{\text{Number of Correct Answers}}{\text{Total Number of Questions}}$
@@ 4.2.2 Baselines With Retrieval
Advanced Retrieval-Augmented Generation: To benchmark ChunkRAG against more sophisticated RAG-based methods, we included advanced models that incorporate additional strategies to enhance performance. SAIL Luo et al. (2023) enhances standard RAG by instruction-tuning the language model on Alpaca instruction-tuning data, inserting top retrieved documents before the instructions to provide contextual information. Self-RAG Asai et al. (2024) further refines RAG by incorporating reflection tokens labeled by GPT-4 within the instruction-tuning data, enabling the model to better utilize retrieved information. Additionally, we considered CRAG Your et al. (2024), a recent approach that augments standard RAG with corrective strategies to improve retrieval quality by addressing low-quality retrieval results. CRAG serves as a direct comparison to our proposed ChunkRAG, highlighting the effectiveness of our chunk filtering mechanism in enhancing retrieval-augmented generation. Furthermore, we evaluated retrieval-augmented baselines with private data, including Ret-ChatGPT and RetLLaMA-chat, which integrate retrieval mechanisms with ChatGPT and the conversational variant of LLaMA2, respectively.